OpenStudy (anonymous):

(Confused!!!!!) What is the arc measure of a sector with area 38.485 ft2 and radius 7?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first use this formula \[\large A = \frac{t}{360}*\pi*r^2\] which is the area of a sector formula

OpenStudy (anonymous):

pi*7^2 = 49*pi =~ 153.938

jimthompson5910 (jim_thompson5910):

you use this formula to determine the angle t once you know the angle t, you can use it to figure out the arc length

OpenStudy (anonymous):

38.485 / (49*pi) =~ 0.25

OpenStudy (anonymous):

so 90 degrees ?

jimthompson5910 (jim_thompson5910):

yeah roughly, nice work so far

jimthompson5910 (jim_thompson5910):

so t = 90 and r = 7 plug them into \[\large s = \frac{t}{360}*2*\pi*r\]

OpenStudy (anonymous):

90/360 * 2 * 3.14 * 7 = I got 175.84

jimthompson5910 (jim_thompson5910):

I'm getting something way smaller

jimthompson5910 (jim_thompson5910):

try it again

jimthompson5910 (jim_thompson5910):

I see where you went wrong, you divided in the wrong order you typed in 360/90 when you should have said 90/360

OpenStudy (anonymous):

10.99 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

:D